OpenStudy (anonymous):

In the function f(x) = 4(x2 − 6x + ____) + 20, what number belongs in the blank to complete the square?

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

The square of half the coefficient of x. (-6/2)² = 9

OpenStudy (anonymous):

so???

OpenStudy (anonymous):

Thats the answer???

OpenStudy (anonymous):

9 is the answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can u show me the work

OpenStudy (anonymous):

i wanna see how u got it

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so to complete the square you take half the coefficient of the 'x' which is 6...so 3...then square it....so 9....and add it to both sides.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

An equation of the form \[ax^2+bx+c=0\] can be written in completed square form as \[a(x+d)^2+e=0\] where \[d=\dfrac{b}{2a}\]\[e=c-\dfrac{b^2}{4a}\]. You have b=-24 and a=4, so d=-24/(2*4)=-24/8=-3. In your question they have the brackets expanded, so we just need to expand the brackets: \[4(x-3)^2+...=4(x^2-6x+9)+...\] (I've omitted the rest as it has no effect on the brackets and you only need to find one number, not the whole lot.